OpenStudy (anonymous):

Find the factors of m2 − 9. (m − 3)(m − 3) (m + 3)(m − 3) (m + 3)(m + 3) Prime

jigglypuff314 (jigglypuff314):

this is difference of two squares: a*a - b*b = (a-b)(a+b) m^2 - 9 = m*m- 3*3

OpenStudy (anonymous):

A?

jigglypuff314 (jigglypuff314):

(a-b)(a+b) ^ ^

OpenStudy (anonymous):

Oh, so B.

jigglypuff314 (jigglypuff314):

yes :)

jigglypuff314 (jigglypuff314):

nice skateboarding videos btw :P

OpenStudy (anonymous):

Haha thanks. I have a passion for film making. :)